The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall Inc. common stock was $ 62 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $ 2,805,100 $ 2,360,600
Net income 668,800 483,500
Total $ 3,473,900 $ 2,844,100
Dividends
On preferred stock $ 8,400 $ 8,400
On common stock 30,600 30,600
Total dividends $ 39,000 $ 39,000
Retained earnings, December 31 $ 3,434,900 $ 2,805,100


Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $ 3,547,800 $ 3,268,770
Cost of goods sold 1,357,800 1,249,180
Gross profit $ 2,190,000 $ 2,019,590
Selling expenses $ 673,700 $ 871,420
Administrative expenses 573,900 511,790
Total operating expenses 1,247,600 1,383,210
Income from operations $ 942,400 $ 636,380
Other income 49,600 40,620
$ 992,000 $ 677,000
Other expense (interest) 232,000 128,000
Income before income tax $ 760,000 $ 549,000
Income tax expense 91,200 65,500
Net income $ 668,800 $ 483,500


Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   Dec. 31, 20Y2    Dec. 31, 20Y1
Assets
Current assets
Cash $ 559,830 $ 750,980
Marketable securities 847,310 1,244,480
Accounts receivable (net) 678,900 635,100
Inventories 511,000 394,200
Prepaid expenses 105,912 150,200
Total current assets $ 2,702,952 $ 3,174,960
Long-term investments 2,593,868 1,697,851
Property, plant, and equipment (net) 3,190,000 2,871,000
Total assets $ 8,486,820 $ 7,743,811
Liabilities
Current liabilities $ 871,920 $ 2,058,711
Long-term liabilities
Mortgage note payable, 8 % $ 1,300,000 $ 0
Bonds payable, 8 % 1,600,000 1,600,000
Total long-term liabilities $ 2,900,000 $ 1,600,000
Total liabilities $ 3,771,920 $ 3,658,711
Stockholders' Equity
Preferred $ 0.70 stock, $ 50 par $ 600,000 $ 600,000
Common stock, $ 10 par 680,000 680,000
Retained earnings 3,434,900 2,805,100
Total stockholders' equity $ 4,714,900 $ 4,085,100
Total liabilities and stockholders' equity $ 8,486,820 $ 7,743,811

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

1. Working capital $ 1831032
2. Current ratio 3.1
3. Quick ratio 2.4
4. Accounts receivable turnover 5.4
5. Number of days' sales in receivables 67.6 days
6. Inventory turnover 3
7. Number of days' sales in inventory 121.7 days
8. Ratio of fixed assets to long-term liabilities 1.1
9. Ratio of liabilities to stockholders' equity 0.8
10. Times interest earned 4.3
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1

11. Asset turnover = Net Sales / Total Assets

= 3,547,800/8,486,820 = 41.8 %

12. Return on total assets = Net Income/Average assets

= 668,800/( 8,486,820+7743811)/2 = 8.2%

13. Return on stockholders' equity = Net Income/Stockholders' Equity

= 668,800/4,714,900 = 14.2%

14. Return on common stock holder equity = Net income - Preference Dividend /Average common stock equity

= 668,800/(4,114,900+3,485,100)/2 = 17.6%

15.Earning per share on common stock = Net Income - Preference dividend/Avg no of shares

= 668,800-8,400/68,000 = $9.71

16.Price-earning ratio =  Price per share/Earning per share

= 62/9.71 = 6.4

17.Dividend per share on common stock= Dividend paid /No of shares

= 30,600/68,000 = $ 0.45

18. Dividend Yield = Annual dividend per share / stock price per share

= 0.45/62 = 0.7%